OpenStudy (mitu12):

find the value of each value. [2x+1 -1] [w-4 9y ] = [-5x-6 5a] [3w+4 9y]

OpenStudy (mitu12):

@DangerousJesse

OpenStudy (anonymous):

is it all one problem ?

OpenStudy (mitu12):

yes i think you figure out each box by its self

OpenStudy (anonymous):

1; 2x 2; w-36y 3;-5x-30a 4;3w+36y thats for each box. i dont know how you would put it as one problem tho?

OpenStudy (mitu12):

2x+w-36y= -5x-30+3w+36y

OpenStudy (anonymous):

x=2w/7+72y/7-30/7 w=7x/2+15--36y y=7x/72+5/12-1/36w

OpenStudy (mitu12):

so do i solve that or is that the answer

OpenStudy (anonymous):

those are the answers

OpenStudy (mitu12):

ohhh thank you so much

OpenStudy (anonymous):

welcomee